OpenStudy (anonymous):

I'm really confused on how to solve this. Anyone able to help? sin4x = -2sin2x

OpenStudy (anonymous):

\[\sin 4x=-2\sin 2x\] \[2\sin 2x \cos 2x+2\sin 2x=0\] \[2\sin 2x \left( \cos 2x+1 \right)=0\] \[Either~\sin 2x=0=\sin n \pi,2x=n \pi,x=n \frac{ \pi }{ 2 }\] or\[\cos 2x=-1=\cos \left( 2n+1 \right)\pi,x=\left( 2n+1 \right)\frac{ \pi }{ 2 }\] here n is an integer.

OpenStudy (anonymous):

I have to do it between 0 & 2pi, so it would be \[\frac{ \pi }{ 2 }\]